# UseRawInputEncoding

### v5

@@ I.7 Transition back to broken symmetric phase
Before the phase transition, a field $\varphi$ rolling down the plateau potential $V=+\frac{1}{2}m^{2}e^{-c\varphi}$ is known to admit an inflation phase with a scale invariant power spectrum lyth . Immediately after phase transition, we continue to have $H_{E}\sim 0$ and at small $m^{2}$, a slow roll inflationary phase in the broken symmetric potential $V_{E}=-\frac{1}{2}m^{2}e^{-c\varphi}+\lambda$, with $c$ moves away from $\infty$ and $\varphi\rightarrow 0$ in the negative $\varphi$ ($\varphi<0$) region.
As $c\varphi$ moves away from $\infty$- ∞, $\phi=e^{c\varphi/2}>0$ rolls away from the phase transition point at $\phi\sim 0$. The renormalised mass squared $m^{2}$ increases and will reach a larger but nearly fixed value $M^{2}$ outside the scaling region. The large mass no longer suppresses the $e^{-c\varphi}$ term. The negative potential $-\frac{1}{2}M^{2}e^{-c\varphi}$ takes on a large negative value which leads to a fast $\varphi$ field roll ($\dot{\varphi}>0$) down the negative potential. As the field speed converts to radiation and matter density (reheats) and $e^{-c\varphi}\rightarrow O(1)$ is reached, the Hubble constant $H_{E}$ is dominated by radiation and matter density. Field roll in this section of the model provides a similar scenario to the standard inflation and hot big bang model except that in this case a slow roll inflationary phase already occurred in the symmetric phase before the phase transition. This will allow for a shorter inflationary phase after phase transition and more importantly, could provide the sought after ”a large patch of smooth background Initial condition” of the inflationary paradigm. The graceful exit could come from the renormalised mass increase due to the system moving away from the scaling region.
Lastly, we consider the large time and large $\varphi>0$ asymptotic solution. This large $\varphi$ limit is the conformal fixed point which Khuory Steinhardt starts its pre-Ekpyrotic roll. One question is whether this same initial condition can be reached again. The field acceleration equation, for $\dot{\varphi}>0$ is
@@
$c\varphi=\ln\phi^{2};\>\>\>\varphi\frac{dc}{d\rho}=\frac{-q}{\rho}.$ (59)
@@
However, we only know that $\lambda$ goes to zero logarithmically in a second order phase transition at small field $\phi$. Since a detailed study of the large field infra-red limit of $\lambda\phi^{4}$ in its broken symmetric phase is not found, it is unclear whether $\lambda$ will reach zero or a finite value in its infra-red limit. There are also factors such as finite size effects due to small spatial curvature and phase transition back to symmetric phase at $m^{2}>0$ that will prevent $\lambda$ goes to zero.
@@
$H_{E}=+\frac{1}{\sqrt{3}}\sqrt{\frac{1}{2}\dot{\varphi}^{2}-\frac{1}{2}m^{2}e^{-c\varphi}+\lambda}=\sqrt{\frac{\lambda}{3}}.$ (61)
Recall Eq.(44), as the $m^{2}/\phi^{2}\rightarrow 0$
@@
If $\lambda_{,\rho}>0$, as $\rho\rightarrow 0$, we have $\dot{\rho}<0$, the last term in Eq.(62) is negative, so that $\dot{H}_{E}<0$. Also from Eq.(58) and Eq.(61), we have $\lambda\rightarrow 0$ and $H_{E}\rightarrow 0$ asymptotically. This is a conformal fixed point.
Alternatively, if we have $\lambda>0$ in the large field infra-red limit, the energy density has a minimum value such that $\rho\rightarrow\rho_{min}$, ($\rho_{min}>0$) and at the point where $\dot{\rho}\rightarrow 0$, we have
@@
We recall that from Eq.(3), we have $\epsilon=\frac{c^{2}}{2}$ and the conformal fixed point corresponds to $c\rightarrow\infty$. When $\dot{\rho}\rightarrow 0$ due to a finite $\lambda>0$, one could still have $\dot{\epsilon}>0$. The end result is that $H_{E}$ takes on a negative value
$H_{E}\sim-\sqrt{\frac{\lambda}{3}}.$ (64)
We have now a contracting universe $H_{E}<0$ and $\dot{H}_{E}<0$, and the energy density will start increasing with $\dot{\rho}>0$.
We consider the field speed $\dot{\varphi}$ in more detail. When the system is close to the asymptotic limit, both $H_{E}$ and $\dot{\varphi}$ take on very small value, but $c$ becomes very large ($(c\sim O(10^{24})$ according to KS) and Eq.(57) will be dominated by the last term as long as $c$ remains large. As $\rho$ reduces, due to the deceleration in Eq.(57), the field speed $\dot{\varphi}>0$ is slowed down to $\dot{\varphi}\sim 0$ and subsequently moves into the reversed speed $\dot{\varphi}<0$ region. In this case, we recover the initial condition for the pre-ekpyrotic phase of KS khoury -khoury1 . One has a cyclical model as postulated by Ijja and Steinhardt steinhardt2
@@ I.8 A Conformal Cyclic Cosmology scenario
Apart from a bounce to a contracting phase, after $\dot{\varphi}<0$, the potential $V_{E}=-\frac{1}{2}m^{2}e^{-c\varphi}+\lambda$ could transition back to $V_{E}=+\frac{1}{2}m^{2}e^{-c\varphi}+\lambda$ in the small $\rho$ region following the Kibble argument kibble -kibble2 . In Jordan frame, the field $\phi$ starts a slow field roll back towards $\phi\sim 0$ region in the symmetric potential. Following this pathway, one could recover a cyclical model without a contracting phase, which shares some features with the conformal cyclic cosmology (CCC) proposed by Penrose penrose . Here the conformal factor $\Omega$ in CCC is identified to $\phi$ in our model as ($\Omega=\phi^{2}$). In CCC, the universe in far future goes into a phase that is cold and the space-time metric is dominated by the Weyl tensor. Here we identify this phase with the conformal fixed point of the field $\phi$. Both models postulate a mechanism which connects the cold phase of a space-time region (aeon) to a hot phase which should describe the observed early universe. Our model here relies only on the conventional phase transition results.
@@ I.9 Summary and Discussion
After Planck (2013), single field plateau potentials beome the simplest choices to provide the inflationary scenario. From a lifted plateau potential, Khoury and Steinhard (KS) argue for an Ekpyrotic solution which could provide a reasonable curvature perturbation spectrum. The challenge is in finding a non-singular bounce mechanism to bring the contracting universe back to the expanding universe. We take the KS potential in Einstein frame to a double well potential in Jordan frame and identify their model constants $V_{0}$, $V_{1}$ with the quartic coupling and mass squared, with $c^{-2}=(K-6)$ measures the deviation of the non-conformal coupling $K$ from its conformal limit. The initial condition for the KS pre-Ekpyrotic phase can be idenfied with the neighbourhood of a conformal fixed point in the Jordan frame system.
@@
In Einstein frame, after the bounce, the scalar field in this expanding universe is expected to roll down a positive plateau potential and reaches the value of quartic coupling $\lambda$ asymptotically. However, in Jordan frame, if the symmetric system is inside the scaling region of a phase transition (which leads to a broken symmetric phase) at small field $\phi$, $\lambda$ matches the value of its broken symmetric phase counterpart which gives the correct present value for the cosmological constant.
@@
We see that once the field $\phi$ rolls to value far away from the conformal point, the renormalised mass increases, the Einstein frame negative potential becomes large (and negative), the field speed $\dot{\varphi}$ from rolling down the potential increases and can be converted to radiation and matter density. Since $c$ reduces rapidly once the system moves away from the conformal point and when $\varphi$ moves across $\varphi=0$ into the positive value region, the negative potential shrinks rapidly in magnitude, the Hubble parameter is dominated by radiation and matter density and the standard big bang scenario is ensued. At very large time (large $\varphi$) and low energy density, the Hubble constant is dominated by the cosmological constant $\lambda$, which can tend to zero at the infra-red conformal fixed point. If $\lambda$ has a ”finite” infra-red limit, one can obtain a bounce (from expansion to contraction) solution away from this infra-red conformal point, which is the same as the initial pre-Ekpyrotic phase condition postulated by KS. At small energy density, if the Jordan frame system transits back to the symmetric phase at large field, it will slow roll back to $\phi\sim 0$ and restart the process. This pathway has no contracting phase and phenomenologically resembles the conformal cyclic cosmology proposed by Penrose.
This model possesses both an inflationary and an ekpyrotic solution, either of which could describe the observed history of our universe. This model could support a cyclical universe paradigm proposed by Ijja and Steinhardt. In each cycle, the scale factor contraction is much smaller than the scale factor increase and therefore the entropy accumulation problem is alleviated. The origin of the universe is therefore pushed back to the first cycles. In our model, there is a minimum scale factor $a_{p}$. In this case, a consistent postulate is that initially the system with a scale factor $a=a_{p}$ sits at a conformally phase transition point, with $\phi=0$, $\dot{\phi}=0$, $K=6$, $m^{2}=\lambda=0$ ($\rho=0$). Because $a\geq a_{p}$, this fixed point is unstable with respect to the broken symmetric phase, so that the system should start by moving into the broken symmetric phase with $a>a_{p}$.

### v6

@@ I.7 Transition back to broken symmetric phase
Before the phase transition, a field $\varphi$ rolling down the plateau potential $V=+\frac{1}{2}m^{2}e^{-c\varphi}$ is known to admit an inflation phase with a scale invariant power spectrum lyth . Immediately after phase transition, from lyth with $t>0$, we continue to have (at small $m^{2}$) a slow roll inflationary phase in the broken symmetric potential $V_{E}=-\frac{1}{2}m^{2}e^{-c\varphi}+\lambda$, with $c$ moves away from $\infty$ and $\varphi\rightarrow 0$ in the negative $\varphi$ ($\varphi<0$) region.
Initially, $e^{-c\varphi}$ is large, but it can be sufficiently suppressed by the small $m^{2}$ term. As $c\varphi$ moves away from $\infty$- ∞, $\phi^{2}=e^{c\varphi}>0$ rolls away from the conformal phase transition point at $\phi\sim 0$. Consequently, the renormalised mass squared $m^{2}$ also moves away from its small value and will increase to a larger but nearly fixed value $M^{2}$ outside the scaling region. This large mass $M$ no longer suppresses the $e^{-c\varphi}$ term. The negative potential $-\frac{1}{2}M^{2}e^{-c\varphi}$ takes on a large negative value which leads to a fast $\varphi$ field roll ($\dot{\varphi}>0$) down the negative potential. As the field speed converts to radiation and matter density (reheats) and $e^{-c\varphi}\rightarrow O(1)$ is reached, the Hubble constant $H_{E}$ is dominated by radiation and matter density. Field roll in this section of the model potential provides a similar scenario to the standard inflation and hot big bang model except that in this case a slow roll inflationary phase already occurred in the symmetric phase before the phase transition. A larger $m^{2}$ that does not full suppress the $e^{-c\varphi}$ term will allow for a shorter inflationary phase after phase transition. Most importantly, a long inflationary era in the late symmetric phase could provide the sought after ”a large patch of smooth background Initial condition” of the inflationary paradigm. The graceful exit comes from the renormalised mass increase due to the system moving away from the scaling region.
Lastly, we consider the large time and large $\varphi>0$ asymptotic solution. This large $\varphi$ limit is the conformal fixed point which Khoury Steinhardt khoury -khoury1 starts its pre-Ekpyrotic roll. One question is whether this same initial condition can be reached again. The field acceleration equation, for $\dot{\varphi}>0$ is
@@
$c\varphi=\ln\phi^{2};\>\>\>c=\frac{1}{\varphi}(\ln\phi_{0}^{2}-q\ln\rho),\>\>\>\>\varphi\frac{dc}{d\rho}=\frac{-q}{\rho}.$ (59)
@@
the R.H.S of Eq.(57) remains negative and in the small $\dot{\varphi}$ region, we have
$\ddot{\varphi}\sim-\frac{q}{\varphi}(\ln\rho)\rho^{l+q}.$ (61)
We consider the case that the small field speed is ”reversed” ($\dot{\varphi}<0$) at some small $\rho$ value. Since $\ddot{\varphi}$ is small in the small $\rho$ region, the field speed $\dot{\varphi}^{2}$ remains small.
@@
$H_{E}=+\frac{1}{\sqrt{3}}\sqrt{\frac{1}{2}\dot{\varphi}^{2}-\frac{1}{2}m^{2}e^{-c\varphi}+\lambda}\sim\frac{1}{\sqrt{3}}\sqrt{\frac{1}{2}\dot{\varphi}^{2}+\lambda}.$ (62)
We recall that from Eq.(28), as the $m^{2}/\phi^{2}\rightarrow 0$
@@
If $\lambda_{,\rho}>0$, as $\rho\rightarrow 0$, we have $\dot{\rho}<0$, the last term in Eq.(63) is negative, so that $\dot{H}_{E}<0$. Also from Eq.(58) and Eq.(62), we have $\lambda\rightarrow 0$ and if $H_{E}\rightarrow\frac{1}{2}\dot{\varphi}^{2}$ asymptotically. This is close to the conformal fixed point at large $\varphi$.
Since we have $\dot{\varphi}^{2}>0$ in the (large field) infra-red limit, the energy density has a minimum value such that $\rho\rightarrow\rho_{min}$, ($\rho_{min}>0$) and at the region where $\dot{\rho}\rightarrow 0$, we have from Eq.(35)
@@
We recall that from Eq.(3), we have $\epsilon=\frac{c^{2}}{2}$ and the conformal fixed point corresponds to $c\rightarrow\infty$, so that
$\frac{\dot{\epsilon}}{\epsilon}=2\frac{\dot{c}}{c}=\frac{\dot{\rho}}{\rho}\frac{2}{\ln\rho}$ (65)
$H_{E}=-\frac{\dot{\rho}}{\rho}\bigg{(}1+\frac{2}{\ln\rho}\ln\bigg{(}\frac{a}{a_{p}}\bigg{)}\bigg{)}$ (66)
For some small $\rho$, from Eq,(66) $H_{E}$ will become negative. The end result is that $H_{E}$ takes on a negative value
$H_{E}\sim-\frac{1}{\sqrt{3}}\sqrt{\frac{1}{2}\dot{\varphi}^{2}-\frac{1}{2}m^{2}e^{-c\varphi}+\lambda}.$ (67)
We now enter a contracting universe $H_{E}<0$ and $\dot{H}_{E}<0$, and the energy density will start increasing with $\dot{\rho}>0$. In this case, we recover the initial condition for the pre-ekpyrotic phase of KS khoury -khoury1 , however one still has to switch to $t<0$ to start the contraction process. One obtains a cyclical model as postulated by Ijja and Steinhardt steinhardt2
@@ I.8 A Conformal Cyclic Cosmology scenario
Apart from a bounce from an expanding phase to a contracting phase, after $\dot{\varphi}<0$, the potential $V_{E}=-\frac{1}{2}m^{2}e^{-c\varphi}+\lambda$ could transition back to $V_{E}=+\frac{1}{2}m^{2}e^{-c\varphi}+\lambda$ in the small $\rho$ region following the Kibble argument kibble -kibble2 . In Jordan frame, the field $\phi$ in the symmetric phase starts a slow field roll back towards $\phi\sim 0$ region. Following this pathway, one could recover a cyclical model without a contracting phase, which shares some features with the conformal cyclic cosmology (CCC) proposed by Penrose penrose . Here the conformal factor $\Omega$ in CCC is identified to $\phi$ in our model as ($\Omega=\phi^{2}$). In CCC, the universe in far future goes into a phase which is inflationary and cold where the space-time metric is dominated by the conformal metric. Here we identify this phase in CCC with the conformal fixed point of the Jordan frame system Eq.(9) with potential in Eq. (21). In our model, to reach the conformal fixed point, both the known massive particles density and the mass of the scalar field (which identifies with the inflaton in an inflationary paradigm) need to be negligible. The non-zero mass of the scalar inflaton arises for phenomenological reasons. Both models postulate the existence of a mechanism which connects the cold phase of a space-time region (aeon) to a hot phase which should describe the observed early universe. CCC postulates that where the universe loses all scales, the cold conformal universe can be matched mathematically to a hot universe. Our model here relies on the conventional phase transition theory and after phase transition, the inflationary phase in the symmetric phase is connected to a hot phase only when the mass term is large enough such that it does not suppress the field term $e^{-c\varphi}$.
@@ I.9 Summary and Discussion
After Planck (2013), single field plateau potentials beome the simplest choices to provide the inflationary scenario. From a lifted plateau potential, Khoury and Steinhard (KS) argue for an Ekpyrotic solution which could provide a reasonable curvature perturbation spectrum. This is a new paradigm and the challenge is in finding a non-singular bounce mechanism to bring the contracting universe back to the expanding universe. We take the KS potential in Einstein frame to a double well potential in Jordan frame and identify their model constants $V_{0}$, $V_{1}$ with the quartic coupling and mass squared, with $c^{-2}=(K-6)$ measures the deviation of the non-conformal coupling $K$ from its conformal limit. The initial condition for the KS pre-Ekpyrotic phase can be idenfied with the neighbourhood of a conformal fixed point in the Jordan frame system.
@@
In Einstein frame, after the bounce, the scalar field in this expanding universe is expected to roll down a positive plateau potential $V_{E}=\frac{1}{2}\frac{m2}{\phi^{2}}+\lambda$ and reaches the value of quartic coupling $\lambda$ asymptotically. If the system is in a phase where the mass term still dominates, one would obtain the correct order of magnitude estimate of the cosmological constant. If the system is in a phase where $\lambda$ dominates, if the symmetric system is inside the scaling region of a phase transition (which leads to a broken symmetric phase) at small field $\phi$, $\lambda$ matches the value of its broken symmetric phase counterpart which gives the correct present value estimate for the cosmological constant.
@@
We see that once the field $\phi$ rolls to value far away from the conformal point, the renormalised mass moves away from its scaling region and increases. The Einstein frame negative potential becomes large (and negative), the field speed $\dot{\varphi}$ from rolling down the potential increases and can be converted to radiation and matter density. Since $c$ reduces rapidly once the system moves away from the conformal point and when $\varphi$ moves across $\varphi=0$ into the positive value region, the negative potential shrinks rapidly in magnitude, the Hubble parameter is dominated by radiation and matter density and the standard big bang scenario is ensued. At very large time (large $\varphi$) and low energy density, the field speed $\dot{\varphi}$ reduces to very small value. We consider the scenario that the field speed decelerate enough to reach a negative field speed in the small field speed region. As energy density reduces, the rate of change in the equation of state parameter $\epsilon$ will lead to a bounce from an expansion phase $H_{E}>0$ to a contraction phase $H_{E}<0$, while $\dot{H}_{E}<0$ remains. This can be identified with the initial pre-Ekpyrotic phase condition postulated by KS . At small energy density, the Jordan frame system can also transit back to the symmetric phase at large field, the large $\phi$ will slow roll back to $\phi\sim 0$ and restart the process. This pathway has no contracting phase and phenomenologically resembles the conformal cyclic cosmology proposed by Penrose.
This model possesses both an inflationary and an ekpyrotic solution, either of which could describe the observed history of our universe. However, the field roll in the symmetric phase at low energy density seems to be a critical component of any realisic model. This model could support a cyclical universe with a bounce phase proposed by Ijja and Steinhardt. It can also support a cyclical universe with no contracting phase which is phenomenologically similar to the CCC proposed by Penrose. The origin of the universe is therefore pushed back to the first cycles. In our model, there is a minimum scale factor $a_{p}$. In this case, a consistent postulate is that initially the system with a minimum scale factor $a=a_{p}$ sits at a conformally invariant phase transition point, with the effective field $\phi=0$, $\dot{\phi}=0$, $K=6$, $m^{2}=\lambda=0$ ($\rho=0$). The field $\phi$ dynamics near a conformal point can be sensitive to other plausible quantum effects such as the Coleman-Weinberg mechanism and will not be pursued here.